OpenStudy (anonymous):

which point is a solution to the system of inequalities y2x-3

OpenStudy (anonymous):

Which point is a solution to the system ? (3, 3) (2, -1) (-1, -2) (-4, 1)

OpenStudy (anonymous):

these are answers

OpenStudy (anonymous):

also need help anther question will give medal

OpenStudy (anonymous):

question 2

OpenStudy (anonymous):

Which of the following systems correctly describe the system of inequalities graphed below? graph of a system of two linear inequalities, first line is a solid line that passes through points (6, 0), (9, 2), and beyond, second line is a dotted line passes through (5, 3), (6, 3), and beyond, graph is shaded in the region that includes point (2, 8)

OpenStudy (anonymous):

this is the graph

OpenStudy (anonymous):

these are multiple choice answers A y>3 Y>=2/3x-4

OpenStudy (anonymous):

b y<3 y>=2/3x-4

OpenStudy (anonymous):

c y>3 y<=2/3-4

OpenStudy (anonymous):

d y<3 y<=2/3-4

OpenStudy (michele_laino):

the shaded region, is given by the intersection of the subsequent regions: 1) a first region which is given by all points whose \(y-\) coordinate is greater than \(3\) 2) a second region which is given by all points have the \(y-\) coordinate which is greater than the corresponding \(y-\) coordinates of the slanted line, so, we have to write the equation of such slanted line first

OpenStudy (michele_laino):

oops.. by all points whose \(y-\) coordinate is greater than the corresponding \(y-\)coordinate of the slanted line...

OpenStudy (michele_laino):

now, the slanted line passes at points \((0,-4)\) and \((6,0)\), so its equation is: \[\frac{{y - {y_1}}}{{{y_2} - {y_1}}} = \frac{{x - {x_1}}}{{{x_2} - {x_1}}}\] where \((x_1,y_1)=(0,-4)\) and \((x_2,y_2)=(6,0)\) Please substitute those coordinates into my equation, and write the equation of slanted line, so we can continue to write the solution of the exercise

OpenStudy (anonymous):

so then 0,4 6,0

OpenStudy (anonymous):

is this the first question or second one

OpenStudy (michele_laino):

no, please they are \((0,-4)\) and \((6,0)\) it is the second one

OpenStudy (anonymous):

oh ok sorry

OpenStudy (michele_laino):

hint: here is next step: \[\frac{{y - \left( { - 4} \right)}}{{0 - \left( { - 4} \right)}} = \frac{{x - 0}}{{6 - 0}}\] please simplify

OpenStudy (michele_laino):

another step: \[\frac{{y + 4}}{4} = \frac{x}{6}\] please continue, now it is easy

OpenStudy (anonymous):

1( y2 + x3 )2

OpenStudy (michele_laino):

If I multiply both sides by \(4\) I get: \[y + 4 = \frac{{4x}}{6}\] and then: \[y + 4 = \frac{{2x}}{3}\] so the equation of slanted line is: \[y = \frac{{2x}}{3} - 4\] am I right?

OpenStudy (anonymous):

x6+y

OpenStudy (anonymous):

oh yea i messed up o my multiplying

OpenStudy (michele_laino):

so the region of the exercise, is the region of all points whose \(y-\) coordinate is greater than \(3\) and it is greater than \(2x/3 - 4\) So, what is the right option?

OpenStudy (michele_laino):

oops... I meant is greater or equal than \(2x/3 -4\)

OpenStudy (michele_laino):

hint: if \(a\) is greater than \(b\) I write \(a >b\)

OpenStudy (anonymous):

it would leave d or c

OpenStudy (michele_laino):

I'm sorry, they are both wrong options

OpenStudy (anonymous):

those are the only ones with greater than or equal sign

OpenStudy (anonymous):

likemyou wrote

OpenStudy (michele_laino):

you have to search for this symbols \(>\) and \( \geqslant \)

OpenStudy (michele_laino):

please try

OpenStudy (anonymous):

yeah thats what this means <= or >=

OpenStudy (anonymous):

it a or c

OpenStudy (anonymous):

its c then @Michele_Laino

OpenStudy (michele_laino):

as I wrote before, C is a wrong option

OpenStudy (anonymous):

i meant to say a

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

sorry i am dumb thank you for your patcience @Michele_Laino

OpenStudy (daisyduck04):

hi and it was wrong it was b

OpenStudy (daisyduck04):

thanks for helping though

OpenStudy (michele_laino):

thanks!! @Cruznatalie150 option B, is a wrong option @daisyduck04

OpenStudy (daisyduck04):

no i am actually talking about the question she helped me with

OpenStudy (michele_laino):

ok! :) @daisyduck04

OpenStudy (anonymous):

before i go @Michele_Laino how would i solve the first one

OpenStudy (anonymous):

just tell me my steps u dnt have to help me you done enough already

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (daisyduck04):

hi your in k12?

OpenStudy (daisyduck04):

why did you block me

OpenStudy (michele_laino):

for example, point \((3,3)\) is not the right option, since its coordinates don't check the inequality \(x<y\)

OpenStudy (anonymous):

did you just substitute them to check

OpenStudy (michele_laino):

in order to solve such exercise, you have to search for point, whose coordinates verify contemporarily both the inequalities

OpenStudy (michele_laino):

correct! @Cruznatalie150 we have to substitute into both inequalities

OpenStudy (anonymous):

so then @Michele_Laino

OpenStudy (michele_laino):

let's consider the second point \((2,-1)\), such coordinates verify the first inequality \(y<x\)?

OpenStudy (anonymous):

i tried it work because -1<2 right so then it would be that one because i tried others

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

now we have to substitute into the second inequality, please try

OpenStudy (anonymous):

oh wait thats right and i did and it didnt work

OpenStudy (anonymous):

-1>1 thats nt true

OpenStudy (michele_laino):

correct! Then let's consider the third point and please repeat the same procedure

OpenStudy (anonymous):

it works @Michele_Laino for the first one because-2<-1 but i got stuck in the 2 equation

OpenStudy (anonymous):

-2>2(-1)-3

OpenStudy (anonymous):

aand got -2.-1

OpenStudy (michele_laino):

for second inequality, we have: left side y= -2 right side = \(2 \cdot (-1)-3=-2-3=-5\) so what can you conlcude?

OpenStudy (anonymous):

then i tried the 4 one it didnt work eaither

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

its c because a negative times negative is a positive right

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

correct! It is option C